OpenStudy (anonymous):

please help...THE SUM OF THREE CONSECUTIVE INTERGERS IS 54. FIND THE NUMBERS

OpenStudy (unklerhaukus):

the the smallest of the consecutive number be \(n\) this makes the three consecutive integers \(n,n+1,n+2\) so the sum of them will be \[(n)+(n+1)+(n+2)\] and we know the sum is \(54\) so we have \[(n)+(n+1)+(n+2)=54\] simplify this and solve for \(n\)

OpenStudy (unklerhaukus):

the the* let the

OpenStudy (anonymous):

THanks you so much! I knew this type of formula would be somewhere lettme figure it out and i'll get back to you

OpenStudy (unklerhaukus):

sure

OpenStudy (anonymous):

n+n+1+n+2=54 3n=54-1-2 3n=51 n= 17

OpenStudy (anonymous):

thanks so much!

OpenStudy (anonymous):

54divided by 3 is 18

OpenStudy (anonymous):

so by simply dividing, it wouldn't be the right answer, so u have to get a letter, and then get the formula

OpenStudy (anonymous):

thanks you, I have a math exam, and this really helped

OpenStudy (unklerhaukus):

you have found \(n=17\) and we set the consecutive integers to be \(n,n+1,n+2\) substitute \(n\) and you'll have the numbers

OpenStudy (anonymous):

huh? oh i thought i was done...ok

OpenStudy (anonymous):

17+18+19

OpenStudy (anonymous):

I didn't know i had to do that...consecutive means...one after the other not the smallest

OpenStudy (anonymous):

not the same*** not the smallest

OpenStudy (unklerhaukus):

yeah consecutive does mean one after the other finally you can check you solution does \(17+18+19=54?\)

OpenStudy (anonymous):

ok thank you